OpenStudy (anonymous):

Let A and B be subsets of U with n(A)=8, n(B)=15, n(A')=12, and n(AnB)=6. Find n(AUB')

OpenStudy (anonymous):

probably helps to draw a venn diagram n(AUB') = n(A) + n(B') - n(AnB') n(AnB') = 8 - 6 = 2